# EmbodiedSAM: Online Segment Any 3D Thing in Real Time

### v1

@@ Abstract
Embodied tasks require the agent to fully understand 3D scenes simultaneously with its exploration, so an online, real-time, fine-grained and highly-generalized 3D perception model is desperately needed. Since high-quality 3D data is limited, directly training such a model in 3D is almost infeasible. Meanwhile, vision foundation models (VFM) has revolutionized the field of 2D computer vision with superior performance, which makes the use of VFM to assist embodied 3D perception a promising direction. However, most existing VFM-assisted 3D perception methods are either offline or too slow that cannot be applied in practical embodied tasks. In this paper, we aim to leverage Segment Anything Model (SAM) for real-time 3D instance segmentation in an online setting. This is a challenging problem since future frames are not available in the input streaming RGB-D video, and an instance may be observed in several frames so object matching between frames is required. To address these challenges, we first propose a geometric-aware query lifting module to represent the 2D masks generated by SAM by 3D-aware queries, which is then iteratively refined by a dual-level query decoder. In this way, the 2D masks are transferred to fine-grained shapes on 3D point clouds. Benefit from the query representation for 3D masks, we can compute the similarity matrix between the 3D masks from different views by efficient matrix operation, which enables real-time inference. Experiments on ScanNet, ScanNet200, SceneNN and 3RScan show our method achieves leading performance even compared with offline methods. Our method also demonstrates great generalization ability in several zero-shot dataset transferring experiments and show great potential in open-vocabulary and data-efficient setting. Code and demo are available at HERE, with only one RTX 3090 GPU required for training and evaluation.
@@ 1 Introduction
Embodied tasks, like robotic manipulation and navigation [20; 3; 36; 35], require the agent to understand the 3D scene, reason about human instructions and make decisions with self-action. Among the pipeline, embodied visual perception is the foundation for various downstream tasks. In embodied scenarios, we hope the 3D perception model to be: (1) online. The input data is a streaming RGB-D video rather than a pre-collected one and visual perception should be performed synchronously with data collection; (2) real-time. High inference speed is needed; (3) fine-grained. It should recognize almost any object appeared in the scene; (4) highly-generalized. One model can be applied to different kinds of scenes and be compatible with different sensor parameters like camera intrinsics. As high-quality 3D data is limited, training such a model in pure 3D is almost infeasible.
@@
Inspired by the great achievements of large language models (LLMs) [38; 4; 1], a series of vision foundation models (VFMs) such as SAM [12] and SEEM [40], have emerged. VFMs are revolutionizing the field of 2D computer vision by their fine-grained, accurate and generalizable segmentation on image pixels. However, less studies have been conducted on developing VFMs for the 3D domain. Since there is much less high-quality annotated 3D data compared with 2D counterparts, it holds great promise to explore the adaptation or extension of existing 2D VFMs for embodied 3D perception. Recently, there are some works [33; 34; 17] that adopt SAM to automatically generate masks on multi-view images of a 3D scene and merge the masks in 3D with projection and iterative merging. While these approaches achieve fine-grained 3D instance segmentation with high generalization ability, they still face some serious problems that hinder their application: (1) they apply SAM on individual images and directly project the 2D masks to 3D point clouds with camera parameters. So the predictions are not geometric-aware, which may produce inconsistent results across different views; (2) they merge per-frame mask predictions in 3D with hand-crafted strategy. E.g., computing geometric similarity between all pairs of masks and merge them according to a threshold, which is inaccurate and very slow; (3) most of them are offline methods based on pre-collected RGB-D frames with 3D reconstruction.
@@
In this paper, we propose a VFM-assisted 3D instance segmentation framework namely EmbodiedSAM (ESAM), which exploits the power of SAM to online segment anything in 3D scenes with high accuracy, fast speed and strong generalization ability. As shown in Figure 1, different from previous 3D SAM methods [33; 31; 34] that project 2D masks to 3D and merge them with hand-crafted strategies, ESAM lifts 2D masks to 3D queries and predicts temporal and geometric-consistent 3D masks with iterative query refinement. Benefit from the 3D query representation, ESAM is also able to fastly merge 3D masks in different frames with simple matrix operations. Specifically, we extract point-wise features from the point clouds projected from depth image. Then we regard the 2D masks generated by SAM as superpoints, which is used to guide mask-wise aggregation by our proposed geometric-aware pooling module, generating 3D queries with one-to-one correspondence to SAM masks. We further present a dual-level query decoder to iteratively refine the 3D queries, which makes the queries efficiently attend with superpoint-wise features and generate fine-grained point-wise masks. Since each 3D instance mask is associated with a query, we can compute similarity between newly predicted 3D masks and previous ones by efficient matrix multiplication in parallel and accurately merge them. To enhance the discriminative ability of query features, we design three representative auxiliary tasks for estimation of geometric, contrastive and semantic similarities. We conduct extensive experiments on ScanNet, ScanNet200, SceneNN and 3RScan datasets. Compared with previous VFM-assisted 3D instance segmentation methods, we improve the accuracy and speed by a large margin while still remain strong generalization ability. Moreover, ESAM can be easily extended to open-vocabulary segmentation. It also shows great potential in data-efficient setting when trained with limited data.
@@
Figure 1: Different from previous 3D SAM methods [33; 31; 34] that project 2D masks to 3D and merge them with hand-crafted strategies, ESAM lifts 2D masks to 3D queries and iteratively refine them to predict accurate 3D masks. With 3D queries, ESAM is also able to fastly merge 3D masks in different frames with simple matrix operations. Take SAM3D [33] for comparison, our ESAM surpasses its performance by 15.7% AP with a more than 20×20$\times$20 × faster speed.
@@ 2 Related Work
VFM-assisted 3D Scene Segmentation: In 2D realm, vision foundation models (VFM) [23; 12; 15] have exploded in growth. Benefit from the large amount of annotated visual data, the 2D VFM shows great accuracy and very strong generalization ability, which makes them work well in zero-shot scenarios. Since there is much less high-quality annotated data in the field of 3D vision than the 2D counterpart, using 2D VFM to assist 3D scene perception becomes a promising direction [25; 33; 31; 34]. UnScene3D [25] considers 2D self-supervised features from DINO [23] to generate initial pseudo masks, which is then iteratively refined with self-training. SAM3D [33] adopts SAM [12] to generate 2D instance masks, which are then projected to 3D space by depth and camera parameters and merged according to the geometry. SAMPro3D [31] maps 3D points in scenes to multi-view 2D images as 3D prompts, which are used to align the 2D masks generated by SAM and cluster the 3D points into instance masks. SAI3D [34] generates 3D primitives on the reconstructed 3D mesh. Then it adopts semantic-SAM to acquire 2D masks with semantic scores, which are connected with the 3D primitives and merged by a graph-based region growing strategy. Our approach also utilizes SAM to assist 3D instance segmentation. Differently, we makes the process of 2D-to-3D projection and 3D mask merging learnable and online. In this way, our ESAM is able to predict more accurate 3D masks and be applied in practical real-time online tasks.
@@
Online 3D Scene Perception: In persuit of embodied AI, real world applications like robotic navigation [3; 36] and manipulation [20] have received increasing attention. Online 3D scene perception, which precisely understands the surrounding 3D scenes from streaming RGB-D videos, becomes the visual basis of these robotic tasks. Early online 3D perception methods process 2D images separately and project the predictions to 3D point clouds, which is followed by a fusion step to merge the predictions from different frames [18; 21]. However, the predictions on 2D image is not geometric and temporal-aware, which makes the fusion step difficult and inaccurate. Fusion-aware 3D-Conv [37] and SVCNN [11] construct data structures to maintain the information of previous frames and conduct point-based 3D aggregation to fuse the 3D features for semantic segmentation. INS-CONV [16] extends sparse convolution [9; 5] to incremental CNN to efficiently extract global 3D features for semantic and instance segmentation. In order to simplify the design of online 3D perception model and leverage the power of the advanced offline 3D architectures, MemAda [32] proposes a new paradigm for online 3D scene perception, which empowers offline model with online perception ability by multimodal memory-based adapters. Different from the previous works, our ESAM lifts SAM-generated 2D masks to accurate 3D mask and corresponding queries, which enables us to efficiently merge the per-frame predictions with high accuracy.
@@ 3.1 Query Lifting and Refinement
Geometric-aware Query Lifting. As SAM does not leverage the information from previous frames, nor does it exploit 3D information from the depth image, directly project $M^{2d}$ to $P$ results in inaccurate and temporal-inconsistent 3D masks. Instead, we aim to lift each 2D mask to a 3D query feature, which enables us to further refine the queries for 3D instance mask generation. Since the 2D binary mask is less informative, here we propose to extract point cloud features from the scene and then regard the 2D masks as indexs to cluster point clouds into superpoints, where queries can be simply selected from the superpoint features. Assuming the point clouds $P\in\mathbb{R}^{N\times 3}$ and there are $M$ masks in $M^{2d}$, we first map $M^{2d}$ to $P$ according to the color-depth correspondence to get superpoint index $S\in\mathbb{Z}^{N}$, where each element in $S$ falls in $[0,M)$. Then we feed $P$ to a 3D sparse U-Net [5] with memory-based adapter [32] to extract temporal-aware 3D features $F_{P}\in\mathbb{R}^{N\times C}$. With $F_{P}$ and $S$, we can pool the point-wise features to superpoint features $F_{S}\in\mathbb{R}^{M\times C}$.
@@
$\boldsymbol{z}^{global}={\rm Agg}(\boldsymbol{z}^{local})\in\mathbb{R}^{C},\ \boldsymbol{z}^{local}={\rm MLP}(\mathcal{P}_{i})\in\mathbb{R}^{|\mathcal{P}_{i}|\times C}$ (1)
@@
$w_{j}={\rm Sigmoid}({\rm MLP}(\boldsymbol{z}_{j}))\in\mathbb{R}^{(0,1)},\ \boldsymbol{z}_{j}=[\boldsymbol{z}^{local}_{j},\boldsymbol{z}^{global}]$ (2)
@@
$F_{S}^{i}=\mathcal{G}(F_{P}^{i})+\boldsymbol{z}^{global},\ \mathcal{G}(F_{P}^{i})={\rm mean}(F_{P}^{i}*[w_{1},...,w_{|\mathcal{P}_{i}|}])$ (3)
Note we enhance the pooled superpoint feature with $\boldsymbol{z}^{global}$ to fully combine the shape-level geometric feature and scene-level 3D U-Net feature. The computation for each superpoint can be parallelized with point-wise MLP and Scatter function [8], so this geometric-aware pooling is actually efficient.
@@
$\hat{Q}_{l}={\rm Softmax}(\frac{\boldsymbol{Q}\cdot\boldsymbol{K}^{T}}{\sqrt{C}}+A_{l})\cdot\boldsymbol{V},\ A_{l}(i,j)=\begin{cases}0&{\rm if}\ M_{l}^{cur}(i,j)\ ={\rm True}\\ -\infty&{\rm otherwise}\end{cases},\ l=0,1,2$ (4)
where $\cdot$ indicates matrix multiplication, $\boldsymbol{Q}$ is the linear projection of $Q_{l}$, $\boldsymbol{K}$ and $\boldsymbol{V}$ are the linear projection of the scene representations $F$. $F$ can be point-wise features $F_{P}$ or superpoint-wise features $F_{S}$. $A_{l}$ is the attention mask derived from the predicted 3D instance masks $M_{l}^{cur}$ in the $l$-th decoder layer. $(i,j)$ indicates $i$-th query attending to $j$-th point or superpoint. Then we feed $\hat{Q}_{l}$ to self-attention layer and feed forward network to get $Q_{l+1}$, followed by a mask generation module to predict the instance mask for each query:
@@
A common practice [27; 28; 14] for qeury decoder is to adopt the same level of scene representations for cross-attention and mask generation. However, since SAM has already outputs high-level semantic-aware masks, we observe $M\ll N$. If we adopt the point-wise scene representations $F_{P}$ for query decoder, the cross-attention operation will be memory-consuming due to the large amount of points. While if we use superpoint features $F_{S}$, the predicted 3D instance masks will only be the combination of superpoints and thus cannot be refined to finer granularity. To get the best of both worlds, our query decoder is designed to be dual-level. For cross-attention in Eq (4), we set $F=F_{S}$ to achieve efficient interaction. While for mask prediction in Eq (5), we set $F=F_{P}$ for fine-grained mask generation. To support masked attention, we pool point mask to superpoint mask before Eq (4):

### v3

@@ Abstract
Embodied tasks require the agent to fully understand 3D scenes simultaneously with its exploration, so an online, real-time, fine-grained and highly-generalized 3D perception model is desperately needed. Since high-quality 3D data is limited, directly training such a model in 3D is infeasible. Meanwhile, vision foundation models (VFM) has revolutionized the field of 2D computer vision with superior performance, which makes the use of VFM to assist embodied 3D perception a promising direction. However, most existing VFM-assisted 3D perception methods are either offline or too slow that cannot be applied in practical embodied tasks. In this paper, we aim to leverage Segment Anything Model (SAM) for real-time 3D instance segmentation in an online setting. This is a challenging problem since future frames are not available in the input streaming RGB-D video, and an instance may be observed in several frames so efficient object matching between frames is required. To address these challenges, we first propose a geometric-aware query lifting module to represent the 2D masks generated by SAM by 3D-aware queries, which is then iteratively refined by a dual-level query decoder. In this way, the 2D masks are transferred to fine-grained shapes on 3D point clouds. Benefit from the query representation for 3D masks, we can compute the similarity matrix between the 3D masks from different views by efficient matrix operation, which enables real-time inference. Experiments on ScanNet, ScanNet200, SceneNN and 3RScan show our method achieves state-of-the-art performance among online 3D perception models, even outperforming offline VFM-assisted 3D instance segmentation methods by a large margin. Our method also demonstrates great generalization ability in several zero-shot dataset transferring experiments and show great potential in data-efficient setting. Code is available [^1] Project page: https://xuxw98.github.io/ESAM/.
@@ 1 Introduction
Embodied tasks, like robotic manipulation and navigation Mousavian et al. (2019); Chaplot et al. (2020); Zhang et al. (2023), require the agent to understand the 3D scene, reason about human instructions and make decisions with self-action. Among the pipeline, embodied visual perception is the foundation for various downstream tasks. In embodied scenarios, we hope the 3D perception model to be: (1) online. The input data is a streaming RGB-D video rather than a pre-collected one and visual perception should be performed synchronously with data collection; (2) real-time. High inference speed is needed for robot planning and control; (3) fine-grained. It should recognize almost any object appeared in the scene; (4) highly-generalized. One model can be applied to different kinds of scenes and be compatible with different sensor parameters like camera intrinsics. As high-quality 3D data is limited, training such a model in pure 3D is almost infeasible.
@@
Inspired by the great achievements of large language models (LLMs) Zhang et al. (2022); Chowdhery et al. (2023); Achiam et al. (2023), a series of vision foundation models (VFMs) such as SAM Kirillov et al. (2023) and SEEM Zou et al. (2023), have emerged. VFMs are revolutionizing the field of 2D computer vision by their fine-grained, accurate and generalizable segmentation on image pixels. However, less studies have been conducted on developing VFMs for the 3D domain. Since there is much less high-quality annotated 3D data compared with 2D counterparts, it holds great promise to explore the adaptation or extension of existing 2D VFMs for embodied 3D perception. Recently, there are some works Yang et al. (2023); Yin et al. (2024); Lu et al. (2023) that adopt SAM to automatically generate masks on multi-view images of a 3D scene and merge the masks in 3D with projection and iterative merging. While these approaches achieve fine-grained 3D instance segmentation with high generalization ability, they still face some serious problems that hinder their application: (1) they apply SAM on individual images and directly project the 2D masks to 3D point clouds with camera parameters. So the predictions are not geometric-aware, which may produce inconsistent results across different views; (2) they merge per-frame mask predictions in 3D with hand-crafted strategy. E.g., computing geometric similarity between all pairs of masks and merge them according to a threshold, which is inaccurate and very slow; (3) most of them are offline methods based on pre-collected RGB-D frames with 3D reconstruction, requiring full data collection before perception.
@@
In this paper, we propose a VFM-assisted 3D instance segmentation framework namely EmbodiedSAM (ESAM), which exploits the power of SAM to online segment anything in 3D scenes with high accuracy, fast speed and strong generalization ability. As shown in Figure 1, different from previous 3D SAM methods Yang et al. (2023); Xu et al. (2023); Yin et al. (2024) that project 2D masks to 3D and merge them with hand-crafted strategies, ESAM lifts 2D masks to 3D queries and predicts temporal and geometric-consistent 3D masks with iterative query refinement. Benefit from the 3D query representation, ESAM is also able to fastly merge 3D masks in different frames with simple matrix operations. Specifically, we extract point-wise features from the point clouds projected from depth image. Then we regard the 2D masks generated by SAM as superpoints, which is used to guide mask-wise aggregation by our proposed geometric-aware pooling module, generating 3D queries with one-to-one correspondence to SAM masks. We further present a dual-level query decoder to iteratively refine the 3D queries, which makes the queries efficiently attend with superpoint-wise features and generate fine-grained point-wise masks. Since each 3D instance mask is associated with a query, we can compute similarity between newly predicted 3D masks and previous ones by efficient matrix multiplication in parallel and accurately merge them. To enhance the discriminative ability of query features, we design three representative auxiliary tasks for estimation of geometric, contrastive and semantic similarities. We conduct extensive experiments on ScanNet, ScanNet200, SceneNN and 3RScan datasets. ESAM achieves both leading accuracy and speed among online 3D perception models. Compared with offline VFM-assisted 3D instance segmentation methods, we improve the performance by a large margin while still remain strong generalization ability. Moreover, ESAM also shows great potential in data-efficient setting when trained with limited data.
@@
Figure 1: Different from previous 3D SAM methods Yang et al. (2023); Xu et al. (2023); Yin et al. (2024) that project 2D masks to 3D and merge them with hand-crafted strategies, ESAM lifts 2D masks to 3D queries and iteratively refine them to predict accurate 3D masks. With 3D queries, ESAM is also able to fastly merge 3D masks in different frames with simple matrix operations. Take SAM3D Yang et al. (2023) for comparison, our ESAM surpasses its performance by 23.2% AP with a more than 20×20$\times$20 × faster speed.
@@ 2 Related Work
VFM-assisted 3D Scene Segmentation: In 2D realm, vision foundation models (VFM) Oquab et al. (2023); Kirillov et al. (2023); Li et al. (2023) have exploded in growth. Benefit from the large amount of annotated visual data, the 2D VFM shows great accuracy and very strong generalization ability, which makes them work well in zero-shot scenarios. Since there is much less high-quality annotated data in the field of 3D vision than the 2D counterpart, using 2D VFM to assist 3D scene perception becomes a promising direction Rozenberszki et al. (2024); Yang et al. (2023); Yin et al. (2024). UnScene3D Rozenberszki et al. (2024) considers 2D self-supervised features from DINO Oquab et al. (2023) to generate initial pseudo masks, which is then iteratively refined with self-training. SAM3D Yang et al. (2023) adopts SAM Kirillov et al. (2023) to generate 2D instance masks, which are then projected to 3D space by depth and camera parameters and merged according to the geometry. SAI3D Yin et al. (2024) generates 3D primitives on the reconstructed 3D mesh. Then it adopts semantic-SAM to acquire 2D masks with semantic scores, which are connected with the 3D primitives and merged by a graph-based region growing strategy. Recently, there are some works Qin et al. (2024); Ye et al. (2024) leverage VFM to segment 3D scenes represented by 3D gaussian, which provide a new perspective for 3D instance segmentation. Our approach also utilizes SAM to assist 3D instance segmentation. Differently, we makes the process of 2D-to-3D projection and 3D mask merging learnable and online. In this way, our ESAM is able to predict more accurate 3D masks and be applied in practical real-time online tasks.
@@
Online 3D Scene Perception: In persuit of embodied AI, real world applications like robotic navigation Chaplot et al. (2020); Zhang et al. (2023) and manipulation Mousavian et al. (2019) have received increasing attention. Online 3D scene perception, which precisely understands the surrounding 3D scenes from streaming RGB-D videos, becomes the visual basis of these robotic tasks. Early online 3D perception methods process 2D images separately and project the predictions to 3D point clouds, which is followed by a fusion step to merge the predictions from different frames McCormac et al. (2017); Narita et al. (2019). However, the predictions on 2D image is not geometric and temporal-aware, which makes the fusion step difficult and inaccurate. Fusion-aware 3D-Conv Zhang et al. (2020) and SVCNN Huang et al. (2021) construct data structures to maintain the information of previous frames and conduct point-based 3D aggregation to fuse the 3D features for semantic segmentation. INS-CONV Liu et al. (2022) extends sparse convolution Graham et al. (2018); Choy et al. (2019) to incremental CNN to efficiently extract global 3D features for semantic and instance segmentation. In order to simplify the design of online 3D perception model and leverage the power of the advanced offline 3D architectures, MemAda Xu et al. (2024) proposes a new paradigm for online 3D scene perception, which empowers offline model with online perception ability by multimodal memory-based adapters. Different from the previous works, our ESAM lifts SAM-generated 2D masks to accurate 3D mask and corresponding queries, which enables us to efficiently merge the per-frame predictions with high accuracy.
@@ 3.1 Query Lifting and Refinement
Geometric-aware Query Lifting. As SAM does not leverage the information from previous frames, nor does it exploit 3D information from the depth image, directly project $M^{2d}$ to $P$ results in inaccurate and temporal-inconsistent 3D masks. Instead, we aim to lift each 2D mask to a 3D query feature, which enables us to further refine the queries for 3D instance mask generation. Since the 2D binary mask is less informative, here we propose to extract point cloud features from the scene and then regard the 2D masks as indexs to cluster point clouds into superpoints, where queries can be simply selected from the superpoint features. Assuming the point clouds $P\in\mathbb{R}^{N\times 3}$ and there are $M$ masks in $M^{2d}$, we first map $M^{2d}$ to $P$ according to the color-depth correspondence to get superpoint index $S\in\mathbb{Z}^{N}$, where each element in $S$ falls in $[0,M)$. Then we feed $P$ to a 3D sparse U-Net Choy et al. (2019) with memory-based adapter Xu et al. (2024) to extract temporal-aware 3D features $F_{P}\in\mathbb{R}^{N\times C}$. With $F_{P}$ and $S$, we can pool the point-wise features to superpoint features $F_{S}\in\mathbb{R}^{M\times C}$.
@@
$\bm{z}^{global}={\rm Agg}(\bm{z}^{local})\in\mathbb{R}^{C},\ \bm{z}^{local}={\rm MLP}(\mathcal{P}_{i})\in\mathbb{R}^{|\mathcal{P}_{i}|\times C}$ (1)
@@
$w_{j}={\rm Sigmoid}({\rm MLP}(\bm{z}_{j}))\in\mathbb{R}^{(0,1)},\ \bm{z}_{j}=[\bm{z}^{local}_{j},\bm{z}^{global}]$ (2)
@@
$F_{S}^{i}=\mathcal{G}(F_{P}^{i})+\bm{z}^{global},\ \mathcal{G}(F_{P}^{i})={\rm mean}(F_{P}^{i}*[w_{1},...,w_{|\mathcal{P}_{i}|}])$ (3)
Note we enhance the pooled superpoint feature with $\bm{z}^{global}$ to fully combine the shape-level geometric feature and scene-level 3D U-Net feature. The computation for each superpoint can be parallelized with point-wise MLP and Scatter function Fey , so this geometric-aware pooling is actually efficient.
@@
$\hat{Q}_{l}={\rm Softmax}(\frac{\bm{Q}\cdot\bm{K}^{T}}{\sqrt{C}}+A_{l})\cdot\bm{V},\ A_{l}(i,j)=\begin{cases}0&{\rm if}\ M_{l}^{cur}(i,j)\ ={\rm True}\\ -\infty&{\rm otherwise}\end{cases},\ l=0,1,2$ (4)
where $\cdot$ indicates matrix multiplication, $\bm{Q}$ is the linear projection of $Q_{l}$, $\bm{K}$ and $\bm{V}$ are the linear projection of the scene representations $F$. $F$ can be point-wise features $F_{P}$ or superpoint-wise features $F_{S}$. $A_{l}$ is the attention mask derived from the predicted 3D instance masks $M_{l}^{cur}$ in the $l$-th decoder layer. $(i,j)$ indicates $i$-th query attending to $j$-th point or superpoint. Then we feed $\hat{Q}_{l}$ to self-attention layer and feed forward network to get $Q_{l+1}$, followed by a mask generation module to predict the instance mask for each query:
@@
A common practice Schult et al. (2022); Sun et al. (2023); Kolodiazhnyi et al. (2024b) for qeury decoder is to adopt the same level of scene representations for cross-attention and mask generation. However, since SAM has already outputs high-level semantic-aware masks, we observe $M\ll N$. If we adopt the point-wise scene representations $F_{P}$ for query decoder, the cross-attention operation will be memory-consuming due to the large amount of points. While if we use superpoint features $F_{S}$, the predicted 3D instance masks will only be the combination of superpoints and thus cannot be refined to finer granularity. To get the best of both worlds, our query decoder is designed to be dual-level. For cross-attention in Eq (4), we set $F=F_{S}$ to achieve efficient interaction. While for mask prediction in Eq (5), we set $F=F_{P}$ for fine-grained mask generation. To support masked attention, we pool point mask to superpoint mask before Eq (4):